# Bipartite Fluctuations of Critical Fermi Surfaces

### v2

@@ VI Discussion and Outlook
In Sec. IV and Sec. V, we perform our calculations of static structure factors at the critical points (e.g., Eq. (37) and Eq. (41)) within the RPA. We expect these results to hold as long as fluctuation effects are not too strong. The Ioffe-Larkin rule $\Pi^{-1}=\Pi_{f}^{-1}+\Pi_{b}^{-1}$, arising from local gauge constraints, is expected to hold at the level of linear response [78]. Regarding the expression of $\Pi_{f}$, as discussed in Ref. [30, 31, 34, 33, 36, 41], the charge susceptibility of the Fermi surface does not contain any singular contributions. We anticipate that the long-wavelength behavior of $\Pi^{\tau\tau}$ is still dominated by $\Pi_{b}$ in the CFT sector, which has a rigid structure described by Eq. (34). To support this conclusion, one could attempt to derive the full expression of $\Pi_{f}(\mathtt{i}\omega,\bm{k})$ at finite $\omega$ and $\bm{k}$, incorporating higher-order diagrams as done in Refs. [34, 46, 44], albeit these calculations were only carried out under the limit $\bm{k}\rightarrow 0$ for conductivity. Such an endeavor proves to be technically challenging, and we defer it to future studies.
The entanglement entropy in the CFL phase at $\nu=1/2$ has been numerically investigated in Ref. [94, 95]. Both of their results exhibit the $L\log(L)$ scaling, resembling that of free fermions, although their overall coefficients differ by a factor of two. This clearly deviates from our result of a boundary law $\mathcal{F}^{c}\sim L$ for bipartite charge fluctuations in Sec. IV.2.1. As we have seen in the gapless Mott insulator, although the charge fluctuations scale as $\mathcal{F}^{c}\sim L$, the spinon Fermi surface still gives rise to the spin fluctuations $\mathcal{F}^{s}\sim L\log(L)$. Therefore, it would be interesting to investigate the bipartite fluctuations of other quantities in the CFL phase that potentially identify the scaling of its entanglement entropy.
@@
[…]
@@
We have not considered the scenario of transitions between Landau FLs and charge-density-wave insulators coexisting with neutral Fermi surfaces [54, 56]. In such cases, the coupling of the density-wave order parameter $O$ to the Fermi surface of $f$-fermions leads to a different Landau damping term $\int_{k}|\omega||O(\omega,\mathbf{k})|^{2}$, which becomes relevant if the scaling dimension satisfies $\Delta[O]<1$. This scenario is likely true, and there is no dynamical decoupling of the boson and fermion sectors at criticality. It is possible that the Landau damping eventually drives the transition to become weakly first-order. Alternatively, there could be a new fixed point for the charge sector with a dynamical exponent $z>1$, as a technically similar theory has been considered in Ref. [96]. At this stage, we lack the theoretical tools to definitively determine the nature of these transitions, so the study of bipartite fluctuations across these exotic quantum critical points will need to be addressed in future research.
This paper focuses on critical Fermi surfaces in two spatial dimensions. However, continuous Mott transitions between Fermi-liquid metals and spin-liquid insulators in three-dimensional systems can be constructed using similar techniques. For instance, in Ref. [97], the same parton construction as described in Eq. (39) was employed, resulting in a critical theory similar to the two-dimensional case. In this context, the spinon-chargon interaction is found to be marginally irrelevant, and the chargon condensation belongs to the 4D XY universality class. Since the system reaches its upper critical dimension, the scaling forms of various physical quantities exhibit logarithmic corrections. Within the vortex theory framework, the U(1) disorder operator is represented by a Wilson surface. When examining its shape dependence, additional singular geometries arise, including corners, cones, and trihedral vertices. It would be intriguing to explore the universal terms in bipartite fluctuations for critical Fermi surfaces in three dimensions.
In Sec. III, we provide a unified theoretical framework for critical Fermi surfaces by introducing vortices in Landau FLs. A conceptually equivalent but technically distinct approach exists for constructing the vortex theory without using partons. To illustrate this idea, consider the CFL-FL transition. One can start with the non-linear bosonization of spinless FLs [47] (also see App. G for a brief review). Vortex excitations are then introduced in the following theory
Here, $\mathcal{L}_{\textrm{FS}}[\phi]$ is the nonlinear theory introduced in Eq. (118), and $J[\phi]$ represents the U(1) current. The dynamical gauge field $a_{\mu}$ is introduced as a Lagrangian multiplier, enforcing the duality relation $J[\phi]=\frac{\mathtt{i}}{2\pi}\star\textrm{d}\tilde{a}$. By considering the vortex sector as in Eq. (26), one can again realize the CFL-FL transition. The essence lies in the fact that the boson $\phi$ encompasses a large number of low-energy excitations, including modes represented by the partons. The advantage of the theory in Eq. (53) lies in its ability to undergo various generalizations by modifying $J[\phi]$, such as for higher-angular momentum channels, as noted in Ref. [98].
@@
One of our key findings highlights the change in scaling behaviors of bipartite charge fluctuations (or the U(1) disorder operator) across interaction-driven “metal-insulator transitions,” which closely mirrors transitions between different phases of higher-form symmetries. Despite recent progress in understanding certain aspects of the low-energy physics of Fermi surfaces from the perspective of LU(1) anomaly [45], and efforts to formulate Fermi-surface dynamics using nonlinear bosonization via the infinite-dimensional Lie group of canonical transformations [47] (also see App. G), a generalized symmetry principle for the quantum phase transitions considered in this paper remains an open question. Nevertheless, our quantitative result Eq. (16) at the critical point, which identifies a universal corner contribution with a positive universal constant $C_{\rho}$, could be tested through quantum Monte Carlo simulations, potentially aiding in the discovery of UV lattice realizations. Given the recent numerical evidence of the CFL phase in twisted $\textrm{MoTe}_{2}$ [57, 58] and the CFL-FL transition reported in Ref. [71], it would be intriguing to explore the possibility of continuous CFL-FL transitions in related models. Numerically verifying Eq. (16) would provide strong evidence for continuous transitions and offer compelling support for the universal critical theory proposed in Refs. [55, 56].
@@ Acknowledgment
We thank Luca Delacrétaz, Hart Goldman, Chao-Ming Jian, Prashant Kumar, Zlatko Papić, Dam Thanh Son, and Senthil Todadri for related discussions. We thank Cenke Xu for discussions and participation in the early stage of the project, and Meng Cheng for communicating unpublished results. This work was supported in part by the Simons Collaboration on Ultra-Quantum Matter, which is a grant from the Simons Foundation (651440), and the Simons Investigator award (990660).
@@ Appendix A $\textrm{LU}(1)$ Anomaly and Structure Factor
$J^{\tau}(\tau,\bm{x})$ $=\frac{\delta\mathcal{S}}{\delta A_{\tau}}=-\mathtt{i}\int\textrm{d}\theta\rho(\tau,\bm{x},\theta),$ (55)
@@
Here, $A_{\mu}=(A_{\tau},A_{\bm{x}})$ is the ordinary background electromagnetic field. We provide two explanations for the inclusion of $\bm{k}_{F}(\theta)$ in $\mathcal{A}_{\bm{x}}$. (1) We want to turn on a background flux $F_{\bm{x}\bm{k}}$ that ensures the canonical commutation relation $[x_{i},k_{j}]=\mathtt{i}\delta_{ij}$ [100]. (2) It aligns with the semiclassical equation of motion $\dot{\bm{k}}=E$. (One may also check Sec. VI. B of Ref. [45]). The LU(1) Ward identity Eq. (57) leads to
@@
$\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{k})=\frac{1}{(2\pi)^{2}}\int\frac{\textrm{d}\omega}{2\pi}\int_{0}^{2\pi}\textrm{d}\theta\frac{1}{\mathtt{i}\omega-\bm{v}_{F}(\theta)\cdot\bm{k}}\left(\bm{k}\times\frac{\textrm{d}\bm{k}_{F}(\theta)}{\textrm{d}\theta}\right)=-\frac{\textrm{Area}}{(2\pi)^{2}},$ (61)
[…]
@@
$\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{k})=\frac{1}{(2\pi)^{2}}\int\frac{\textrm{d}\omega}{2\pi}\int_{0}^{2\pi}\textrm{d}\theta\frac{k_{F}|\bm{k}|\cos\theta}{\mathtt{i}\omega-v_{F}|\bm{k}|\cos\theta}=\frac{k_{F}|\bm{k}|}{2\pi}\int_{0}^{2\pi}\frac{\textrm{d}\theta}{2\pi}\left(-\frac{|\cos\theta|}{2}\right)=-\frac{k_{F}|\bm{k}|}{2\pi^{2}}.$ (62)
@@
$\hat{\Pi}^{\tau\tau}(\mathtt{i}\omega,\bm{k})=\frac{\Pi^{\tau\tau}(\mathtt{i}\omega,\bm{k})}{1-u\Pi^{\tau\tau}(\mathtt{i}\omega,\bm{k})},$ (63)
Here, $\Pi^{\tau\tau}(\mathtt{i}\omega,\bm{k})$ represents the free-fermion result from Eq. (30), and $u$ denotes a coupling constant. Consequently, the equal-time function becomes
$\int\frac{\textrm{d}\omega}{2\pi}\hat{\Pi}^{\tau\tau}(\mathtt{i}\omega,\bm{k})=-\frac{k_{F}|\bm{k}|}{2\pi^{2}}\frac{\sqrt{2\tilde{u}+1}+2\tilde{u}\arctan(\sqrt{2\tilde{u}+1})}{(2\tilde{u}+1)^{3/2}},$ (64)
where $\tilde{u}=u\mathscr{D}_{F}$ is rescaled by the density of states $\mathscr{D}_{F}=\frac{k_{F}}{2\pi v_{F}}$ at the fermi level. Also, note that in the problem of the half-filled Landau level, the Ward identity (Eq. (57)) needs modification due to the presence of an emergent gauge field [45]. Therefore, the geometric interpretation presented here should not conflict with the results in Sec. IV.2.1.
Appendix B Coulomb Gauge and Response Theory
In this appendix, we introduce our convention about the response theory under the Coulomb gauge. For any rotationally invariant systems, the two-point function $\Pi^{\mu\nu}=\langle J^{\mu}J^{\nu}\rangle$ of conserved currents has the structure

### v3

@@ VI.1 Non-Fermi Liquids
In Sec. IV and Sec. V, we computed the static structure factors at the critical points (e.g., Eq. (40) and Eq. (44)) within the RPA framework. We expect these results to remain valid as long as gauge fluctuations are not too strong. The Ioffe-Larkin rule $\Pi^{-1}=\Pi_{f}^{-1}+\Pi_{b}^{-1}$, which arises from local gauge constraints, is expected to hold at the level of linear response [88]. Regarding the expression for $\Pi_{f}$, prior studies [31, 32, 35, 34, 37, 42] have shown that the charge susceptibility of the Fermi surface does not exhibit any singular contributions. We therefore anticipate that the long-wavelength behavior of $\Pi^{\tau\tau}$ remains dominated by the bosonic contribution $\Pi_{b}$ in the CFT sector, whose structure is rigidly constrained and given by Eq. (37). To further support this conclusion, one could attempt to compute the full expression for $\Pi_{f}$ at finite $\omega$ and $\bm{k}$ by including higher-order diagrams, as partially done in Refs. [35, 47, 45]. However, existing calculations are typically restricted to the $\bm{k}\rightarrow 0$ limit relevant for conductivity. Extending these computations beyond this limit is technically demanding, and we leave such analysis to future work.
We have not considered the transitions between Landau FLs and charge-density-wave insulators, each coexisting with a neutral Fermi surface [55, 57]. In such cases, the coupling of the density-wave order parameter $O$ to the Fermi surface of $f$-fermions leads to a different Landau damping term $\int_{k}|\omega||O(\omega,\mathbf{k})|^{2}$, which becomes relevant if the scaling dimension satisfies $\Delta[O]<1$. This scenario is likely true, and there is no dynamical decoupling of the boson and fermion sectors at criticality. It is possible that the Landau damping eventually drives the transition to become weakly first-order. Alternatively, there could be a new fixed point for the charge sector with a dynamical exponent $z>1$, as explored in a technically similar context in Ref. [107]. At this stage, we lack the theoretical tools to definitively determine the nature of these transitions, so the study of bipartite fluctuations across these exotic quantum critical points will need to be addressed in future research.
While this paper focuses on critical Fermi surfaces in two spatial dimensions, continuous Mott transitions between Fermi-liquid metals and spin-liquid insulators in three dimensions can also be constructed using similar theoretical frameworks. For example, Ref. [108] employs the same parton construction as described in Eq. (42), leading to a critical theory analogous to the two-dimensional case. In three dimensions, the interaction between spinons and chargons is found to be marginally irrelevant, and the chargon condensation falls into the 4D XY universality class. Since the system reaches its upper critical dimension, the scaling forms of various physical quantities exhibit logarithmic corrections. Within the vortex theory framework, the U(1) disorder operator is represented by a Wilson surface. When examining its shape dependence, additional singular geometries arise, including corners, cones, and trihedral vertices. It would be interesting to investigate whether universal subleading terms also appear in the bipartite fluctuations of critical Fermi surfaces in 3+1D systems.
In Sec. III, we present a unified parton framework for critical Fermi surfaces by introducing vortices in Landau FLs. A conceptually equivalent but technically distinct approach exists for constructing the vortex theory without using partons. To illustrate this idea, consider the CFL-FL transition. One may begin with the non-linear bosonization of spinless FLs [48] (also see App. H for a brief review). Vortex excitations are then incorporated through the theory
Here, $\mathcal{L}_{\textrm{FS}}[\phi]$ is the nonlinear theory introduced in Eq. (132), and $J[\phi]$ represents the U(1) current. The dynamical gauge field $a_{\mu}$ appears as a Lagrange multiplier that enforces the duality relation $J[\phi]=\frac{\mathtt{i}}{2\pi}\star\textrm{d}\tilde{a}$. By considering the vortex sector as in Eq. (29), one can again realize the CFL-FL transition. The key insight is that the bosonic field $\phi$ encompasses a large number of low-energy excitations, including modes represented by the partons. A notable advantage of the formulation in Eq. (54) lies in its capacity for systematic generalization by modifying $J[\phi]$, for example to incorporate higher-angular-momentum channels, as mentioned in Ref. [109].
VI.2 Quantum Hall Physics
In Sec. IV, we examine the CFL phase, where the localization tensor takes the isotropic form $\mathsf{G}^{ij}=\mathsf{G}\delta^{ij}$ and a subleading corner contribution is identified in Eq. (12). Based on the standard RPA analysis of the HLR theory, we find that the value of $\mathsf{G}$ is universal and is given by the Hall conductivity $\sigma_{xy}$ (see Eq. (13)). However, as we will clarify in an upcoming work [71], the situation is different for general (anisotropic) charge insulators. In such cases, $\mathsf{G}^{ij}$ becomes non-universal and is constrained by both a topological lower bound and an energetic upper bound [64, 65, 66, 82] (see also App. E)
$0$ $\leq\mathsf{G}^{ij}+\frac{\mathtt{i}}{2}\sigma_{xy}\epsilon^{ij},$
$0$ $\leq\lim_{\bm{A}\rightarrow 0}\frac{1}{2\Delta}\frac{1}{V}\left\langle\!\frac{\partial^{2}H}{\partial A_{i}\partial A_{j}}\!\right\rangle-\mathsf{G}^{ij},$ (55)
where the right-hand side of each line is a positive semi-definite matrix. Here, $\Delta$ denotes the optical gap of the many-body Hamiltonian $H$, $V$ is the volume of the system, and $\bm{A}$ represents the background U(1) field. The specific conditions under which these bounds are saturated will be discussed further in Ref. [71].
Under certain circumstances, the corner contribution to bipartite fluctuations can serve as a diagnostic for QH transitions. For instance, the expression $\mathsf{G}^{ij}=\delta^{ij}|\sigma_{xy}|/2$ is guaranteed by Galilean invariance in systems governed by microscopic Hamiltonians of the form [84, 71]
$H=\sum_{\mathsf{j}}\frac{\bm{\pi}_{\mathsf{j}}^{2}}{2m}+\sum_{\mathsf{i}\neq\mathsf{j}}U(|\bm{r}_{\mathsf{i}}-\bm{r}_{\mathsf{j}}|),$ (56)
where $\bm{\pi}_{\mathsf{j}}=\bm{p}_{\mathsf{j}}-\bm{A}(\bm{r}_{\mathsf{j}})$ is the kinetic momentum of $\mathsf{j}$-th particle with mass $m$, and $U$ is an arbitrary interaction potential depending only on pairwise distances. Any QH state realized by such a UV description satisfies the bipartite fluctuations described by Eq. (12) and Eq. (13). Consequently, changes in the corner term can signal topological phase transitions. In real materials, the bounds in Eq. (55) are typically not saturated. However, if one starts with a gapped QH state and varies the parameters, a sudden drop in the value of $\mathsf{G}^{ij}$ below the topological lower bound would still signal a topological transition. Such a transition—between a fractional Chern insulator and a charge-density-wave state driven by a displacement field—was recently discussed in Ref. [110] in the context of twisted bilayer semiconductors.
@@ VI.2 Quantum Hall Physics
$\mathcal{L}_{\textrm{vrtx}}[\psi,\tilde{a}]=\sum_{I=1}^{4}\bar{\psi}_{I}\not{\textrm{D}}_{\tilde{a}}\psi_{I}+\frac{1}{2e^{2}}\textrm{d}\tilde{a}\wedge\star\textrm{d}\tilde{a}$. (57)
@@
VI.3 Entanglement Properties
The study of the corner contribution to the entanglement entropy in CFTs has a longer history [111, 112, 113, 114]. Considering the shape shown in Fig. 1, the entanglement entropy takes the form $L-\textrm{s}(\theta)\log(L)+\ldots$ where #\# is again non-universal. While the analytical expression for $\textrm{s}(\theta)$ is not known in general, it has been shown that, in the limit where $\theta$ approaches $\pi$, $\textrm{s}(\theta)$ is proportional to $C_{T}(\pi-\theta)^{2}$, where $C_{T}$ is the stress-tensor central charge of the CFT [113, 114]. The universal subleading term has been utilized in Monte Carlo studies of deconfined quantum critical points [115, 116, 117, 27], as well as symmetric mass generation [26]. However, there are subtleties in correctly extracting the subleading term and accounting for contributions from Goldstone modes [118, 119].
The entanglement entropy in the CFL phase at $\nu=1/2$ has been numerically investigated in Ref. [120, 121]. Both of their results exhibit the $L\log(L)$ scaling, resembling that of free fermions, although their overall coefficients differ by a factor of two. This problem was recently revisited by Ref. [100], which provided evidence supporting an enhancement in entanglement compared to free fermions. This result clearly deviates from our result of a boundary law $\mathcal{F}^{c}\sim L$ for bipartite charge fluctuations in Sec. IV.2.1. A similar situation arises in the gapless Mott insulator, where the entanglement entropy has been numerically observed to exhibit a logarithmic enhancement [122]. As we have seen in Sec. V, although the charge fluctuations scale as $\mathcal{F}^{c}\sim L$, the spinon Fermi surface still gives rise to the spin fluctuations $\mathcal{F}^{s}\sim L\log(L)$. Therefore, it would be interesting to investigate the bipartite fluctuations of other quantities in the CFL phase that potentially identify the scaling of its entanglement entropy.
The entanglement entropy in both phases separated by the critical points discussed in Sec. IV and Sec. V scales as $L\log(L)$. It is therefore natural to expect the critical points to follow the same scaling, although further careful technical studies are needed. While critical Fermi surfaces and CFTs share similarities in their bipartite charge fluctuations, they are likely to be distinguished by their entanglement entropy.
In the context of non-interacting fermions, the full entanglement spectrum is entirely encoded in the bipartite fluctuations and the higher-order cumulants [^7] More precisely, the full entanglement spectrum of non-interacting fermions depends only on cumulants with even $n$ [3, 4, 5, 6, 7, 8, 9, 10, 11, 12].
$\mathcal{N}_{\mathcal{A}}^{[n]}=\lim_{\vartheta\rightarrow 0}(-\mathtt{i}\partial_{\vartheta})^{n}\log\langle\mathcal{W}_{\mathcal{A}}(\vartheta)\rangle,$ (58)
where $\mathcal{W}_{\mathcal{A}}(\vartheta)$ denotes the U(1) disorder operator
$\mathcal{W}_{\mathcal{A}}(\vartheta)=\exp\left(\vartheta\int_{\mathcal{A}}\star J\right).$ (59)
In the dual vortex theory, the disorder operator corresponds to the Wilson loop Eq. (24), where $\star J=\frac{\mathtt{i}}{2\pi}\textrm{d}\tilde{a}$. A comprehensive understanding of the higher-order cumulants in critical Fermi surfaces requires a systematic extension of the Ioffe-Larkin rule to gauge-invariant non-linear response functions. It would be interesting to explore whether there are any distinctive features between even and odd values of $n$, as observed in Ref. [123] for integer QH and Laughlin states.
A closely related topic is the role of the density three-point function in probing Fermi-surface topology in free fermion systems [78, 79]. It would be interesting to explore how this result generalizes to the non-Fermi liquids considered in this work. Moreover, this may offer a new perspective for distinguishing critical Fermi surfaces from CFTs, since $\langle JJJ\rangle$ vanishes identically for an abelian conserved spin-1 current $J$ in any 3D CFT [124].
VI.4 Numerics and Experiments
@@
One of our key findings highlights the change in scaling behaviors of bipartite fluctuations (or U(1) disorder operators) across interaction-driven “metal-insulator transitions,” which closely mirrors transitions between different phases of higher-form symmetries. Despite recent progress in understanding certain aspects of the low-energy physics of Fermi surfaces from the perspective of LU(1) anomaly [46], and efforts to formulate Fermi-surface dynamics using nonlinear bosonization via the infinite-dimensional Lie group of canonical transformations [48] (also see App. H), a generalized symmetry principle for the quantum phase transitions considered in this paper remains an open question. Nevertheless, our quantitative results in Eq. (14) and Eq. (16) have direct implications for numerical and experimental studies of unconventional quantum criticalities in metals.
Our analytical result Eq. (16), which identifies a universal corner contribution with a universal constant $C_{\rho}$, provides a numerical method (via the Monte Carlo approach) to self-consistently verify whether a proposed candidate microscopic model realizes the desired critical field theory. In the context of CFTs, this method has been benchmarked using the Bose-Hubbard model [19] and applied to study unconventional phase transitions, such as the symmetric mass generation transition of Dirac fermions [26]. The numerical method based on our approach would be especially valuable for verifying the realizability of intriguing quantum critical points in metals, where powerful techniques such as the conformal bootstrap [68, 69] are not applicable.
Significant progress has recently been made in the search for microscopic realizations of the CFL-FL transition. On the theoretical side, a candidate model based on spatially modulated Landau levels was discussed in Ref. [81], where tuning the interaction strength leads to a transition characterized by the vanishing of the Drude weight. Although whether the transition is continuous remains to be determined, this setup provides a promising platform for exploring critical behaviors. In parallel, experimental signatures consistent with zero-field CFL phases and CFL-FL transitions have been observed in twisted $\textrm{MoTe}_{2}$ bilayers [60, 61] and multilayer graphene systems [62]. These developments offer new opportunities for both numerical and experimental tests of the universal scaling behaviors discussed in this work.
Several experimental techniques have been proposed for measuring bipartite fluctuations. In cold atoms and quantum gases, correlation functions and bipartite fluctuations can be directly probed using high-resolution imaging techniques (see e.g. [125, 126, 127, 128, 129]). For 1+1D systems exhibiting space-time duality [4, 11], charge fluctuations can be obtained through quantum point contact measurements [130, 131].
In quantum materials, the localization tensor $\mathsf{G}$ (i.e., the many-body quantum metric), which governs the corner term in $2+1$ dimensions [71, 72], can also be experimentally accessed. For example, $\mathsf{G}$ can be extracted from structure factor measurements using inelastic X-ray scattering, as demonstrated by recent experimental data for the insulator LiF [132]. Furthermore, it can be obtained through frequency sum rules from optical responses, with a detailed analysis for the $\textrm{MnBi}_{2}\textrm{Te}_{4}$ film provided in Ref. [133]. These techniques suggest feasible pathways for experimentally testing the universal predictions and investigating the broader landscape of unconventional quantum criticalities discussed in this work.
@@ Acknowledgment
We thank Luca Delacrétaz, Dominic Else, Hart Goldman, Chao-Ming Jian, Prashant Kumar, Michael Levin, Zlatko Papić, Dam Thanh Son, and Senthil Todadri for related discussions. We thank Cenke Xu for discussions and participation in the early stage of the project, and Meng Cheng for communicating unpublished results. This work was supported in part by the Simons Collaboration on Ultra-Quantum Matter, which is a grant from the Simons Foundation (651442), and the Simons Investigator award (990660).
@@ Appendix A $\textrm{LU}(1)$ Anomaly and Structure Factor
$J^{\tau}(\tau,\bm{x})$ $=\frac{\delta\mathcal{S}}{\delta A_{\tau}}=-\mathtt{i}\int\textrm{d}\theta\rho(\tau,\bm{x},\theta),$
@@
Here, $A_{\mu}=(A_{\tau},A_{\bm{x}})$ is the ordinary background electromagnetic field. We provide two explanations for the inclusion of $\bm{k}_{F}(\theta)$ in $\mathcal{A}_{\bm{x}}$. (1) We want to turn on a background flux $F_{\bm{x}\bm{k}}$ that ensures the canonical commutation relation $x_{i},k_{j}]=\mathtt{i}\delta_{ij} [$135]. (2) It aligns with the semiclassical equation of motion $\dot{\bm{k}}=E$. (One may also check Sec. VI. B of Ref. [46]). The LU(1) Ward identity Eq. (62) leads to
@@
$\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{k})=\frac{1}{(2\pi)^{2}}\int\frac{\textrm{d}\omega}{2\pi}\int_{0}^{2\pi}\textrm{d}\theta\frac{-1}{\mathtt{i}\omega-\bm{v}_{F}(\theta)\cdot\bm{k}}\left(\bm{k}\times\frac{\textrm{d}\bm{k}_{F}(\theta)}{\textrm{d}\theta}\right)=\frac{\textrm{Area}}{(2\pi)^{2}},$ (66)
where $(\textrm{Area})$ represents the value of the shaded area in Fig. 2. The dependence on the fermi velocity $\bm{v}_{F}$ is eliminated after the $\omega$-integral, and $\bm{k}\times\partial_{\theta}\bm{k}_{F}(\theta)$ has the geometric interpretation of an area element.
@@
$\Pi^{\tau\tau}(\tau\rightarrow 0,\bm{k})=\frac{1}{(2\pi)^{2}}\int\frac{\textrm{d}\omega}{2\pi}\int_{0}^{2\pi}\textrm{d}\theta\frac{-k_{F}|\bm{k}|\cos\theta}{\mathtt{i}\omega-v_{F}|\bm{k}|\cos\theta}=\frac{k_{F}|\bm{k}|}{2\pi}\int_{0}^{2\pi}\frac{\textrm{d}\theta}{2\pi}\frac{|\cos\theta|}{2}=\frac{k_{F}|\bm{k}|}{2\pi^{2}}.$ (67)
@@
$\hat{\Pi}^{\tau\tau}(\mathtt{i}\omega,\bm{k})=\frac{\Pi^{\tau\tau}(\mathtt{i}\omega,\bm{k})}{1+u\Pi^{\tau\tau}(\mathtt{i}\omega,\bm{k})},$ (68)
Here, $\Pi^{\tau\tau}(\mathtt{i}\omega,\bm{k})$ represents the free-fermion result from Eq. (33), and $u$ denotes a coupling constant. Consequently, the equal-time correlation function becomes
$\int\frac{\textrm{d}\omega}{2\pi}\hat{\Pi}^{\tau\tau}(\mathtt{i}\omega,\bm{k})=\frac{k_{F}|\bm{k}|}{2\pi^{2}}\frac{\sqrt{2\tilde{u}+1}+2\tilde{u}\arctan(\sqrt{2\tilde{u}+1})}{(2\tilde{u}+1)^{3/2}},$ (69)
where $\tilde{u}=u\mathscr{D}_{F}$ is rescaled by the density of states $\mathscr{D}_{F}=\frac{k_{F}}{2\pi v_{F}}$ at the fermi level. Also, note that in the problem of the half-filled Landau level, the Ward identity Eq. (62) needs modification due to the presence of an emergent gauge field [46]. Therefore, the geometric interpretation presented here should not conflict with the results in Sec. IV.2.1.
Appendix B Response Theory and Coulomb Gauge
The imaginary-time (Matsubara) response function of a $\mathrm{U}(1)$-conserved current is defined as
$\Pi^{\mu\nu}(x-y)=\left.-\frac{\delta^{2}\log\mathcal{Z}[A]}{\delta A_{\mu}(x)\delta A_{\nu}(y)}\right|_{A=0}=\left.\left\langle\frac{\delta^{2}\mathcal{S}[A]}{\delta A_{\mu}(x)\delta A_{\nu}(y)}\right\rangle\right|_{A=0}-\langle J^{\mu}(x)J^{\nu}(y)\rangle_{c}$ (70)
where $\mathcal{Z}[A]=\int\mathcal{D}[\textrm{matter}]e^{-\mathcal{S}[\textrm{matter},A]}$ is the partition function in the presence of the background field $A_{\mu}$, and
$\langle J^{\mu}(x)J^{\nu}(y)\rangle_{c}=\langle J^{\mu}(x)J^{\nu}(y)\rangle-\langle J^{\mu}(x)\rangle\langle J^{\nu}(y)\rangle$ (71)
denotes the connected two-point function of the paramagnetic current operator, which itself is defined as
$J^{\mu}(x)=\left.\frac{\delta\mathcal{S}[A]}{\delta A_{\mu}(x)}\right|_{A=0}.$ (72)
For nonrelativistic electrons, the diamagnetic contact term is absent in the density channel in both the Landau FL and insulating phases, i.e., $\delta^{2}\mathcal{S}/(\delta A_{\tau}\delta A_{\tau})=0$. By applying the Källén-Lehmann spectral representation, one can show that the static structure factor defined in the real-time formalism coincides with the equal-time limit $\tau\rightarrow 0^{+}$ of the Matsubara density response function.
For any translationally and rotationally invariant systems in 2+1 dimensions, the Fourier transform of the response function, i.e., $\Pi^{\mu\nu}(k)=\int\textrm{d}^{3}xe^{-\mathtt{i}x\cdot k}\Pi^{\mu\nu}(x)$, has the structure